OpenStudy (anonymous):

A cylindrical can is to have the volume of 1 L. Find the height and radius that will minimize the surface area. I did the question but My answer is off.

OpenStudy (anonymous):

\[ V=\pi r^2h=1\implies h=\frac{1}{\pi r^2}\\ \] here is the trick.. Can has a curved surface. It has a circular base. What about the top? would it be open? if open, do not consider that area

OpenStudy (anonymous):

\[S=2\pi rh+\pi r^2+\pi r^2\]

OpenStudy (anonymous):

the top is closed. I did the question using s=2pirh+2pir^2 for surface area. then sub in h= 1000/pir^2 (volume formula that I isolated h for). then I found the derivative